# Applications of Generative AI (GAI) for Mobile and Wireless Networking: A Survey

### v1

@@ VI Challenges and Future Directions
Table VI: Summary of Challenges, Critical Issues, and the Respective Solutions of GAI in Mobile Networking
Challenges Issues Potential Impact on Mobile Networking Applications Future Research Directions
Handling technical complexity Improved performance and cost trade-offs in GAI-based mobile networks Develop resource-efficient GAI models
Technical Challenges Tackling legacy system Integration challenges and security risks Explore SDN and network slicing integration
Avoiding technical debt Improved quality and reliability of GAI models Implement continuous learning and improvement
Potential misuse and AI hallucinations Generation of harmful or biased content Develop content authentication and moderation methods
Data handling violations Protection of personal data and privacy rights Implement data anonymization and encryption
Human-AI interaction Ethical and social implications in mobile network workforce Explore human-in-the-loop approaches
Integration challenges Compatibility issues with existing infrastructure Develop new network paradigms and standards
Scalability Resource limitations for large-scale GAI deployment Explore cloud and edge computing solutions
Complexity Difficulty in interpreting complex GAI models Focus on transparent and interpretable algorithms
Lack of Interpretability Diversity Challenges in comparing and integrating diverse data sources Utilize visualization for diverse data
Uncertainty Difficulty in dealing with probabilistic and stochastic models Employ quantification for uncertainty estimation
Unintended consequences Unforeseen outcomes from GAI decisions Develop interpretability metrics and evaluation
User trust Lack of user trust due to opaque decisions Enhance interpretability for user confidence
Authenticity Risk of fake or misleading content generation Develop digital watermarking and verification methods
Ethical and Legal Concerns Ownership Determining ownership and intellectual property rights Implement digital attribution and licensing systems
Fairness Bias and unfairness in content generation Incorporate fairness mechanisms
Responsibility Accountability and responsibility for GAI-generated outcomes Integrate responsibility by design principles
Societal impact Broader ethical implications on society Address broader ethical considerations in GAI systems
Unauthorized access Breach of system security Implement strong authentication and access controls
Security, Privacy, and Trust Data leakage Unauthorized sharing of sensitive information Enhance data encryption and secure communication
Trustworthiness assessment Difficulty in evaluating GAI reliability Develop trust evaluation frameworks and metrics
Adversarial attacks Targeted attacks on GAI models Enhance security against adversarial attacks
VI-A Technical Challenges
To enable new services and use cases for future mobile networks (i.e., SemCom, mobile edge computing, user behavior analysis, wireless signal processing, channel rendering, 3D mapping, and integrated sensing), GAI faces some technical challenges that need to be addressed. The details of each challenge and the corresponding solution are discussed in the following.
Handling technical complexity: GAI models often require powerful computing ability and data to train and fine-tune, necessitating mobile network components to sustain the trade-off between performance and cost. For instance, ChatGPT-3, a text generation tool, requires 175 billion parameters and hundreds of billions of words from the internet. In this context, exploiting cloud or edge computing services is crucial to reducing computation and storage processing tasks from mobile devices or networks. Also, integrating model compression approaches into GAI systems like pruning, quantization, or distillation are also needed to reduce model size and complexity without affecting performance.
Tackling the legacy system: The incompatibility of GAI models with existing network infrastructure and protocols can limit their integration and scalability. In particular, current mobile networks, developed based on centralized servers and fixed bandwidth resources, may not work well with the GAI model and can even generate risks involving network security, privacy, and reliability. To cope with these issues, there is a need for new network paradigms such as SDN and network slicing to provide more flexibility, efficiency, and customization for GAI applications. In addition to this, it is also necessary to develop new standards and regulations to ensure the interoperability of GAI models on mobile networks.
Avoiding technical debt: The risk of falling behind the latest developments and best practices in GAI can affect the quality and reliability of services. Mobile network operators must keep up with the latest research and innovation in the field and update the deployed AI models to avoid errors, bugs, or vulnerabilities that may arise over time. Following that, two possible solutions should be considered: (1) employing a continuous learning and improvement approach to monitor the performance and feedback of GAI models and to make adjustments accordingly and (2) collaborating with other stakeholders such as researchers, developers, or customers to share general AI knowledge and best practices.
Potential misuse and AI hallucinations: GAI models without control limitations can compromise the quality and reliability of content, potentially producing fake news, spam, phishing emails, and malware. Attackers can exploit these to manipulate users or compromise their security. Hence, feedback and evaluation mechanisms, such as digital watermarking or blockchain, are needed to verify the authenticity of GAI content, along with moderation or filtering techniques to detect harmful content. To further prevent security threats related to the generation of adversarial examples, GAI-based systems must have strong and flexible security measures, such as encryption, authentication, anomaly detection, and prevention. In particular, GAI models may require access to personal data without proper consent or protection, resulting in sensitive information leaks. To ensure data privacy, GAI-based mobile systems need to have mechanisms to encrypt, anonymize, or delete data after use.
Human-AI interaction: The integration of GAI into mobile networks to perform/replace some tasks can impact the skills and roles of the workforce, as well as the ethical and social aspects of employment, education, training, and motivation. This raises questions about GAI models’ accountability, transparency, fairness, and trustworthiness. To address this, human-in-the-loop or human-on-the-loop approaches are needed to create a tight connection between human workers and GAI in terms of design, operation, and evaluation. At the same time, building ethical principles and guidelines is critical to ensure the responsible and beneficial use of GAI in mobile networks as well as the shared values and norms of users and providers.
VI-B Lack of Interpretability
Lack of interpretability, another challenge of GAI for mobile networks, refers to lack of explaining and understanding the reasons behind GAI-based system outputs, particularly important in critical or sensitive tasks like network security, performance optimization, or service personalization. Interpretability acts as a bridge between mobile users and providers, fostering trust and confidence. For example, errors, failures, or unexpected outputs produced by GAI-based systems (i.e., generating harmful or biased content), interpretability becomes a powerful tool to help users and providers debug and improve the quality of GAI models. When there are growing concerns regarding ethical or legal issues (e.g., data violations, copyright infringement, or misinformation), interpretability will be the basis to assist users and providers in solving accountability and responsibility problems. It is not straightforward to obtain these benefits due to the following reasons.
Complexity: GAI-based systems are often developed based on complex and non-linear algorithms. For example, DNN contains a lot of hidden layers, parameters, and interactions that are difficult to interpret or visualize. Hence, GAI-based systems should adopt simpler or more transparent algorithms, such as linear models, decision trees, or rule-based methods.
Diversity: As the inputs of GAI models are diverse and heterogeneous (i.e., text, audio, video, or code), it is hard to perform comparison and integration due to their different formats, structures, and even semantics. A possible solution to overcome such challenges is to exploit visualization features to enhance the diversity of GAI-based systems due to graphical and interactive tools (e.g., charts, diagrams, or dashboards), which allow analysis and visualization of the data or outputs of GAI models more intuitively or understandably.
Uncertainty: Most GAI models rely on probabilistic and stochastic models to produce estimated outputs, such as GANs. Therefore, the randomness, variability, and noise during the training process complicate the measurement, processing, and control. The use of quantification features, in this case, is a reasonable choice to mitigate the uncertainty of GAI-based systems, e.g., numerical or statistical approaches such as metrics, scores, or confidence intervals, as well as in estimating the accuracy and reliability of the outputs.
VI-C Ethical and Legal Concerns
Ethical and legal concerns arise when applying GAI to mobile networks as they relate to the ethical and legal implications of creating, using, and sharing new content. These may affect the rights, interests, values, and responsibilities of network stakeholders, such as developers, users, operators, and regulators; the law; and society in general.
Authenticity: In mobile networking, GAI-generated content introduces concerns regarding the authenticity and reliability of the information. This requires measures like digital watermarking to establish source identity and origin identification as well as digital verification techniques (i.e., cryptography or statistics) to ensure content integrity. More importantly, mobile network users need to be further educated with digital literacy-oriented programs to have a basis for evaluating and verifying GAI-generated content and its sources.
Ownership: The emergence of creative content from GAI systems raises concerns about ownership, authorship, and intellectual property rights in mobile networks. Digital rights management solutions become essential to protect and manage these rights, ensuring fair use and preventing unauthorized distribution. GAI-based systems should incorporate digital attribution mechanisms to acknowledge the contributions of original data sources from wireless networks and mobile devices. Furthermore, the development of digital licensing systems is crucial for specifying and enforcing terms and conditions governing the use, modification, and distribution of GAI-generated content.
Fairness: GAI systems raise concerns about bias in the contents generated from data sets, algorithms, or even with creators. Ensuring fairness characteristics of diversity and inclusion requires long-term investments during the deployment of GAI systems for mobile networks. In that, implementing fairness mechanisms, including auditing, debiasing, and calibration, becomes crucial to detect and rectify biases or errors. In addition to this, establishing and enforcing fairness policies are vital steps in defining rules and obligations for GAI systems in active mobile networking environments.
Responsibility: Impactful information generated by GAI-based systems raises concerns about responsibility and accountability. Determining who is liable for the outcomes of GAI systems and how to ensure adherence to ethical and legal standards becomes paramount, particularly for mobile networking applications. Addressing these concerns is not limited to responsibility mechanisms (e.g., feedback, reporting, and review processes) to monitor and assess the impacts of GAI systems but also involves cooperation between developers and users in establishing policies to uphold ethical norms and regulations in the mobile networking context.
VI-D Security, Privacy, and Trust
Although GAI gains significant enhancements in the performance and efficiency of mobile networks, there are various challenges and risks associated with integrating GAI into mobile networks, especially in terms of security, privacy, and trust. In what follows, we will discuss the challenges and the respective solutions for these issues.
VI-D1 Security
Security refers to the ability to protect GAI-based systems and the data they use or produce from unauthorized access, modification, and damage. In particular, GAI can be vulnerable to cyberattacks, such as hacking, spoofing, DoS, or adversarial learning. These attacks can compromise the integrity, availability, and confidentiality of the network and its data. In particular, when GAI is poisoned, it becomes a powerful tool for attackers to launch attacks on other systems/users or to manipulate/extract sensitive information.
To ensure security in mobile networks with GAI, some measures could be taken. First, there is a need to establish security policies and rules as the basis for operating, updating, and maintaining the safety of GAI-based systems. Second, it is required to implement security mechanisms to protect GAI systems from external or internal attacks, such as encryption, authentication, authorization, firewalling, monitoring, and auditing. Third, there is a need for security testing to evaluate and verify the security performance and resilience of GAI systems under different scenarios and conditions. Finally, security awareness and education are needed to raise the knowledge and skills of GAI users and operators on how to prevent and respond to security incidents.
VI-D2 Privacy
Privacy refers to the ability to respect and protect the personal or sensitive information of users. Specifically, the integration of GAI into mobile networks could pose a risk to user data privacy. In that, GAI could collect, process, store, or share personal or confidential data without proper consent, transparency, or accountability. Moreover, GAI could leak sensitive information from the data, such as identity, location, behavior, preferences, or health status. This could violate the privacy rights and expectations of the network stakeholders.
To cope with the privacy issues in GAI-empowered mobile networks, some potential measures should be considered. First, it is required to use privacy mechanisms to safeguard personal or confidential data, such as anonymization, pseudonymization, encryption, access control, and data minimization. Second, there is a need to establish policies and rules to maintain user privacy in GAI systems. Third, privacy tests are necessary to reduce the potential privacy risks and harms that could be caused by GAI systems. Finally, privacy notices and choices is needed to inform the network stakeholders about their privacy rights and options.
VI-D3 Trust
Trust refers to the ability to ensure the reliability, accuracy, and quality of GAI-based systems and the outputs they generate. Normally, GAI could undermine the perception of reliability, competence, honesty, and benevolence of a system or a provider by being unpredictable, inconsistent, and opaque. For example, GAI could make errors or mistakes that affect the quality or availability of the network services. In addition, GAI could hide or falsify its actions that affect the security or privacy of the network.
To tackle such problems, some potential solutions can be exploited for GAI-based systems. First, there is a need to deploy trust mechanisms for GAI systems and their interactions, such as reputation, feedback, certification, verification, and explanation. Second, policies and rules are needed to maintain the trustworthiness of GAI systems to users. Third, a greater focus on measuring and monitoring trust assessments across diverse criteria and metrics will help increase the reliability and performance of the GAI system. Finally, trust education and communication programs are needed to enhance the credibility of GAI systems and the benefits of stakeholders.

### v2

@@ VI Challenges and Future Directions
Table VII: Summary of Challenges, Respective Solutions, and Future Research Directions of GAI in Mobile and Wireless Networking.
Challenges Description Possible Solution Future Research Directions
Complexity • High training and testing time • Overparameterization of inputs • Lightweight features: Lightweight tuning paradigm [219], distributed GAI networking solutions like innovative generative IoV architecture [220] or online generative FL model [221]. • Effective learning approaches: Finite-difference score matching [222]. Future research in GAI-integrated networks should prioritize developing GAI in forms of distributed networking like generative FL or generative collaboration in FL. Additionally, the design of onboard-integrated tiny/lightweight GAI models is also important for compatibility with IoT devices.
Scalability • Transferring model • Division of infrastructure networking • Customizable attributes: Unsupervised learning BertNet approach coupled with lightweight tuning features [149, 219]. • Flexible network control paradigm: hybrid federated and centralized learning [223], knowledge-defined network model [224], virtual network function [225], and multi-modal GAI paradigm [226]. Future research in GAI-integrated networks should focus on developing GAI modules with customizable options to be compatible with diverse network components. At the same time, there is a need for a designed network control paradigm with GAI plugging functions.
Security, Privacy, and Trust • Heterogeneity of network components • Differential security protocols • Consumer’s perspective on GAI use • Enhancing Data Privacy and Integrity: Adoption of physical countermeasures [110] or defence strategies against the threats of GAI [227] and developing GAI-based supply chain security and blockchain security mechanisms [228]. • Robust Single/Multilayered Protocols: Using an automatic generation of distributed GAI algorithms [229] and defence-in-depth by leveraging layers of controls protected by a hardware-enhanced trusted execution environment [230]. • Building Trustworthy in GAI Interactions: Evaluating AI trustworthiness [231] or transforming ‘black box’ GAI models into more transparent and understandable ‘glass box’ models [232]. Future research in GAI-based networks should continuously improve security and privacy mechanisms at the physical-layer transmission while developing GAI frameworks oriented towards supply chain and blockchain securities. Besides, it is also required to design automatic distributed GAI algorithms combined with multilayered security protocols and realize GAI performance frameworks for evaluating trustworthiness in GAI-based systems.
Standardization • GAI model evaluation • Rule and policy in use, operation, and maintenance • GAI model design: Using a comprehensive GAI evaluation framework [233]. • Standardization bodies: Relying on 3GPP standardization [234, 235, 236] for the 5G core network (OAM and RAN) and enhanced 5G NR air interface [237]. Future research should prioritize creating a general GAI framework, defining the mechanisms needed, and exploring new use cases for GAI. These efforts lay the foundation for 6G, which will integrate GAI and communication as key features.
VI-A Complexity
Integrating GAI into mobile and wireless networks can significantly benefit various advanced applications and services. However, GAI-based LMs typically require large computational resources and abundant data for training and fine-tuning [145, 146, 147]. For instance, ControlNet [96] requires a significant amount of training time from scratch until the model is well performed. Specifically, the model is trained for 300 GPU hours using NVIDIA A100 80GB GPUs, taking user scribbles as input conditions, and for 400 GPU hours using the same GPUs for the semantic segmentation task. Additionally, the pre-trained GAI-based LMs are not always available for specific tasks, causing a significant growth in the training time with the number of task requests. Moreover, compared to traditional DNNs, the inference duration (i.e., testing time) in GAI-based diffusion models is considerably longer due to the progressive denoising process. Evidently, the DDPM [74] requires up to 1,000 iterative steps to denoise a single data sample, with each step taking approximately 10 seconds. Importantly, GAI-based models are often constructed with highly complex architectures. For instance, ChatGPT-3, a text generation tool, requires 175 billion parameters, while StableDiffusion v1.5 [238] uses 1 billion ones (6.04 GB) for both UNet and CLIP, which serve as image and text encoders.
To overcome the aforementioned challenges of IoT infrastructure networks, where IoT components do not always have sufficient capabilities (i.e., size, memory, and computational requirements) to effectively perform GAI-based LM processes (e.g., unmanned aerial vehicles [190]), GAI-integrated framework should be lightweight. In this regard, a lightweight tuning paradigm in [219] can be adopted to minimize the fundamental tasks of labeling data for training GAI models while achieving flexible re-modulation of attention and adaptation of pre-trained weights, making GAI modules more flexible for low-resource scenarios and enhancing their ability to transfer knowledge across IoT domains. In addition, exploiting distributed networking solutions for GAI is also crucial for reducing computation and storage tasks. For example, one can generalize the innovative generative Internet-of-Vehicles (IoV) architecture to establish a collaborative fine-tuning mechanism for pretrained GAI models, along with a self-adaptive harmony search-based resource allocation strategy [220]. This framework allows resource-limited IoT devices to achieve rapid, customized, and lightweight GAI, by optimizing time and energy consumption through collaborative fine-tuning. Another solution is to consider online generative FL, as in [221], to deal with non-independent and identically distributed (non-IID) trajectory data across data owners in travel time estimation applications. Specifically, the framework establishes the generative global model shared by all clients to infer real-time road traffic conditions and adjusts a personalized model for each client to analyze their driving habits, thereby correcting the residual errors from the global model’s localized predictions. Moreover, designing effective learning approaches is also needed, such as the finite-difference score matching approach in [222], for ML applications to parallelly approximate any-order directional derivative, without gradient computation during training. In summary, computational requirements and high complexity hugely affect the feasibility of GAI-based IoT solutions. Thus, it is expected that further efforts will be made in these suggested directions.
VI-B Scalability
GAI is typically trained on specific LMs (e.g., CLIP [94], BLIP [95]) or specific tasks (e.g., ControlNet [96]) with constrained learning properties. The scalability of GAI in transferring one domain to others is quite limited in the context of IoT [148], where IoT data often exhibit varying characteristics. Besides, the division of networking infrastructures into centralized and decentralized components, such as computing architecture, data models, and application protocols, further complicates the seamless integration of GAI into existing mobile and wireless networks [239]. Moreover, to flexibly adapt to varying requirements from service providers, GAI-based frameworks must incorporate high-level features that enable them to specify the desired network capabilities.
From the mentioned above, future works should first focus on developing GAI models with strong learning capabilities, which are less affected by classification tasks, such as the unsupervised learning approach of BertNet in [149], coupled with a lightweight tuning paradigm in [219]. To accommodate both centralized and decentralized networking deployments, developing hybrid GAI models that leverage the advantages of distributed and centralized learning [223] is a promising approach. This means that only clients with sufficient computational power and resources will perform FL, while the clients with limited capability will directly transfer their data to the centralized server for processing centralized training tasks. Besides, to provide intelligent network services with real-time decision-making capabilities, GAI-based frameworks for SDN must have robust data processing and quick response features. For example, applying the knowledge-defined network modelling approach in [224] can build a self-driving SDN environment with better precision and efficient performance under variable traffic loads and delays, even with disparate routing schemes in a flat topology. Meanwhile, using the virtual network function solution in [225], which enables dynamic adaptation and reconfiguration of orchestrator deployments based on demand, can proactively achieve autoscaling functions in a cloud-native network. To achieve high dynamicity in interpreting the user’s intent and ensure compatibility with various orchestrators or new operating systems, exploiting the multi-modal GAI paradigm in [226] is also a promising approach.
VI-C Security, Privacy, and Trust
The inherent diversity and complexity of emerging IoT applications and services (e.g., intelligent transportation, smart homes/cities, healthcare, financial sectors, academia, Industry 5.0, and military aspects like joint radar and communication) have led to several challenges to GAI integration, where any IoT component, from data to devices, can be a potential vulnerability point, including GAI-deployed components. Notably, when a GAI-based security module is poisoned, it can become a powerful tool for attackers, combined with other attacking approaches [110], to manipulate/extract sensitive information without resistance [50]. In addition, each IoT element or group is operated and controlled by different security protocols (e.g., sensor-to-sensor, sensor-to-IoT node, mobile-to-sensor, etc.), presenting unique challenges. These factors introduce diverse security risks and create considerable gaps between security measures. Below are possible strategies for these risks.
Enhancing Data Privacy and Integrity. To enhance user privacy, future works should first focus on constructing countermeasures from the physical-layer transmission perspectives, such as clustering-based methods, differential privacy, encryption and homomorphic encryption, generative steganography strategies, and secure multi-party computation [110]. Moreover, the adoption of defence strategies against the threats of GAI is also needed to enhance the personal or confidential data from the potential privacy loss and risks during data exchange and aggregation, such as red and blue teaming to identify system vulnerabilities, adversarial training to strengthen GAI models, explainable AI to interpret trained model outputs, security awareness and training to address the risks associated with the misuse of GAI, and input/output filtering to identify, mitigate, and potentially reject suspicious results [227]. To ensure the data integrity, it is required to implement security mechanisms such as authentication, authorization, and continuous monitoring. For example, developing GAI-based supply chain security approaches can create digital maps of supply chain routes, locations, and intermediaries for tracking the journey of goods, identify potential vulnerabilities, and exploit blockchain security mechanisms for creating digital signatures or hashes and for monitoring blockchain network activities and transactions to detect unusual patterns or behaviours [228].
Robust Single/Multilayered Protocols. Unlike classical security, the evolving nature of cyber threats continuously adapts to AI variations. Consequently, GAI-secured systems must automate and proactively update their measurement mechanisms to protect against emerging vulnerabilities [240]. For example, an automatic generation of distributed GAI algorithms in [229] can be adopted for automating the process of fault-tolerant distributed algorithms, such as reliable broadcast, total order broadcast, causal broadcast, and consensus. Moreover, developing multi-layered or defense-in-depth strategies is crucial to ensure that if one protective layer is breached, the other layers continue to offer protection. An example of a multi-layered strategy includes a firewall as the first line of defense, an IDS as the second line, and regular vulnerability assessments as the third line. Another example is to apply the principle of defence in depth by leveraging layers of controls protected by a hardware-enhanced trusted execution environment [230].
Building Trustworthy in GAI Interactions. The traditional design of ML and AI models has focused solely on achieving accurate output predictions. However, the recent shift towards generative IoT is evident in the emergence of generative search and recommendation systems capable of content retrieval, reuse, and creation to meet user needs. Given GAI’s significant impacts on science, health, and humanity (i.e., ethical and legal concerns), the development of future intelligent systems must focus on achieving accurate results while ensuring transparency, fairness, and unbiasedness. This consequently raises the following important questions:
  • What does it mean to trust a GAI, and how can we evaluate its trustworthiness?
  • What are the mechanisms/tools for building trustworthy GAI, and what is the role of interpretable GAI in trust?
To what extent these questions significantly enhance trust in GAI-based systems has not yet been answered. However, a similar AI approach is referenced in [231]. This involves first distinguishing human-GAI trust from human-GAI-human trust in conjunction with evaluating a GAI’s contractual trustworthiness. Next, using a ladder of model access helps define the role of interpretability in trust. Finally, creating behaviour certificates that are more accurate, relevant, and understandable, rather than merely opening up black boxes, is essential. Another feasible solution is to transform ‘black box’ GAI models into more transparent and understandable ‘glass box’ models [232]. This approach highlights the clear importance of transparency for stakeholders (i.e., users, developers, and policymakers) as well as the trade-offs between transparency and other objectives, such as accuracy and privacy.
VI-D Standardizations
GAI’s mobile and wireless networking integration is a promising strategy for the development of advanced 5G and 6G networks. Moreover, factors like the complexity in LLM for training, data fine-tuning, incompatibility in infrastructure networks, communication protocols, and control models, as well as ethical and legal issues, hinder GAI applications. Since GAI is in its early stages of evolution for various purposes, standardizing its involved learning model/architecture/commercial modules remains an unanswered question. For example, how can we define lightweight GAI models? Answering this question requires a comprehensive GAI evaluation framework [233], with rigorous sequential steps from dataset constructions, feature extractors (types of learning models), and score calculators, to the methods for model evaluation and enhancement. Another question is how to integrate GAI into mobile and wireless networks and operate it. This is an extreme challenge to standardization bodies (e.g., IEEE, ITU, and 3GPP) and big techs (e.g., KT Corp, Samsung, and Cisco). As one of the first aspirants in this direction, 3GPP has recently engaged in preliminary AI/ML initiatives for the 5G core network, with emphasis on two sectors: the operations, administration, and maintenance (OAM) and the radio access network (RAN), through 3GPP documents TR 28.908, TS 28.105, and TR 37.817 [234, 235, 236]. To clarify the potential of AI/ML-based algorithms in enhancing the 5G New Radio (NR) air interface, 3GPP has recently released the latest version of 3GPP TR 38.843 [237].